OpenStudy (anonymous):

!!!! HELP ME PLEASE !!!! Completing the square: a. 32x^2-6x+1=0 b. x^2-x=-4 c. 2x^2+x=-3

OpenStudy (rockstar0765):

whats the question

OpenStudy (anonymous):

Solve it by completing the square

OpenStudy (rockstar0765):

sorry cant help right now i forgot how to do these

OpenStudy (tkhunny):

Look very, very hard at this: \((a+b)^{2} = a^{2} + 2ab + b^{2}\)

OpenStudy (muzzack):

1 Step by step solution : Step 1 : Simplify 32x2-6x + 1 Trying to factor by splitting the middle term 1.1 Factoring 32x2-6x+1 The first term is, 32x2 its coefficient is 32 . The middle term is, -6x its coefficient is -6 . The last term, "the constant", is +1 Step-1 : Multiply the coefficient of the first term by the constant 32 • 1 = 32 Step-2 : Find two factors of 32 whose sum equals the coefficient of the middle term, which is -6 . -32 + -1 = -33 -16 + -2 = -18 -8 + -4 = -12 -4 + -8 = -12 -2 + -16 = -18 -1 + -32 = -33 1 + 32 = 33 2 + 16 = 18 4 + 8 = 12 8 + 4 = 12 16 + 2 = 18 32 + 1 = 33 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Equation at the end of step 1 : 32x2 - 6x + 1 = 0 Step 2 : Solve 32x2-6x+1 = 0 Parabola, Finding the Vertex : 2.1 Find the Vertex of y = 32x2-6x+1 Parabolas have a highest or a lowest point called the Vertex . Our parabola opens up and accordingly has a lowest point (AKA absolute minimum) . We know this even before plotting "y" because the coefficient of the first term, 32 , is positive (greater than zero). Each parabola has a vertical line of symmetry that passes through its vertex. Because of this symmetry, the line of symmetry would, for example, pass through the midpoint of the two x -intercepts (roots or solutions) of the parabola. That is, if the parabola has indeed two real solutions. Parabolas can model many real life situations, such as the height above ground, of an object thrown upward, after some period of time. The vertex of the parabola can provide us with information, such as the maximum height that object, thrown upwards, can reach. For this reason we want to be able to find the coordinates of the vertex. For any parabola,Ax2+Bx+C,the x -coordinate of the vertex is given by -B/(2A) . In our case the x coordinate is 0.0938 Plugging into the parabola formula 0.0938 for x we can calculate the y -coordinate : y = 32.0 * 0.09 * 0.09 - 6.0 * 0.09 + 1.0 or y = 0.719 Parabola, Graphing Vertex and X-Intercepts : Root plot for : y = 32x2-6x+1 Axis of Symmetry (dashed) {x}={ 0.09} Vertex at {x,y} = { 0.09, 0.72} Function has no real roots Solve Quadratic Equation by Completing The Square 2.2 Solving 32x2-6x+1 = 0 by Completing The Square . Divide both sides of the equation by 32 to have 1 as the coefficient of the first term : x2-(3/16)x+(1/32) = 0 Subtract 1/32 from both side of the equation : x2-(3/16)x = -1/32 Now the clever bit: Take the coefficient of x , which is 3/16 , divide by two, giving 3/32 , and finally square it giving 9/1024 Add 9/1024 to both sides of the equation : On the right hand side we have : -1/32 + 9/1024 The common denominator of the two fractions is 1024 Adding (-32/1024)+(9/1024) gives -23/1024 So adding to both sides we finally get : x2-(3/16)x+(9/1024) = -23/1024 Adding 9/1024 has completed the left hand side into a perfect square : x2-(3/16)x+(9/1024) = (x-(3/32)) • (x-(3/32)) = (x-(3/32))2 Things which are equal to the same thing are also equal to one another. Since x2-(3/16)x+(9/1024) = -23/1024 and x2-(3/16)x+(9/1024) = (x-(3/32))2 then, according to the law of transitivity, (x-(3/32))2 = -23/1024 We'll refer to this Equation as Eq. #2.2.1 The Square Root Principle says that When two things are equal, their square roots are equal. Note that the square root of (x-(3/32))2 is (x-(3/32))2/2 = (x-(3/32))1 = x-(3/32) Now, applying the Square Root Principle to Eq. #2.2.1 we get: x-(3/32) = √ -23/1024 Add 3/32 to both sides to obtain: x = 3/32 + √ -23/1024 In Math, i is called the imaginary unit. It satisfies i2 =-1. Both i and -i are the square roots of -1 Since a square root has two values, one positive and the other negative x2 - (3/16)x + (1/32) = 0 has two solutions: x = 3/32 + √ 23/1024 • i or x = 3/32 - √ 23/1024 • i Note that √ 23/1024 can be written as √ 23 / √ 1024 which is √ 23 / 32 Solve Quadratic Equation using the Quadratic Formula 2.3 Solving 32x2-6x+1 = 0 by the Quadratic Formula . According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = 32 B = -6 C = 1 Accordingly, B2 - 4AC = 36 - 128 = -92 Applying the quadratic formula : 6 ± √ -92 x = ————— 64 In the set of real numbers, negative numbers do not have square roots. A new set of numbers, called complex, was invented so that negative numbers would have a square root. These numbers are written (a+b*i) Both i and -i are the square roots of minus 1 Accordingly,√ -92 = √ 92 • (-1) = √ 92 • √ -1 = ± √ 92 • i Can √ 92 be simplified ? Yes! The prime factorization of 92 is 2•2•23 To be able to remove something from under the radical, there have to be 2 instances of it (because we are taking a square i.e. second root). √ 92 = √ 2•2•23 = ± 2 • √ 23 √ 23 , rounded to 4 decimal digits, is 4.7958 So now we are looking at: x = ( 6 ± 2 • 4.796 i ) / 64 Two imaginary solutions : x =(6+√-92)/64=(3+i√ 23 )/32= 0.0938+0.1499i or: x =(6-√-92)/64=(3-i√ 23 )/32= 0.0938-0.1499i Two solutions were found : x =(6-√-92)/64=(3-i√ 23 )/32= 0.0938-0.1499i x =(6+√-92)/64=(3+i√ 23 )/32= 0.0938+0.1499i

OpenStudy (tkhunny):

See that \(a^{2}\)? Get something square up front! This is first. a. 32x^2-6x+1=0 b. x^2-x=-4 c. 2x^2+x=-3 Line them up right. a. 32x^2 - 6x + ____ = -1 b. x^2 - x + ____ = -4 c. 2x^2 + x + ____ = -3 Get something square in the front. a. 2(16x^2 - 3x + ____) = -1 b. x^2 - x + ____ = -4 c. 2(x^2 + ½x + ____) = -3 Adjust the other side to prepare for the magic number to come. a. 2(16x^2 - 3x + ____) = -1 + 2(____) b. x^2 - x + ____ = -4 + ____ c. 2(x^2 + ½x + ____) = -3 + 2(____) Find something square to put on the end.: a. (3/2)^2 = 9/4 b. (1/2)^2 = 1/4 c. (½/2)^2 = 1/16 Fill in the Blanks a. 2(16x^2 - 3x + 9/4) = -1 + 2(9/4) b. x^2 - x + 1/4 = -4 + 1/4 c. 2(x^2 + ½x + 1/16) = -3 + 2(1/16) Rewrite in the SQUARE version. a. 2(4x - 3/2)^2 = 7/2 b. (x - ½)^2 = -4 c. 2(x + ¼)^2 = -23/8 That was fun. You do the rest.

OpenStudy (anonymous):

Oh my god thnks so much guys!!!

OpenStudy (tkhunny):

I don't actually recommend doing three problems at the same time. :-)